Calculate the divergence of the following vector functions:

Vb = xy X +2yzY +3zx Z


1
Expert's answer
2021-03-10T17:14:58-0500

"div \vec{V_b}=\frac{\partial V_{bx}}{\partial x}+\frac{\partial V_{by}}{\partial y}+\frac{\partial V_{bz}}{\partial z}=y+2z+3x" . Answer